Im really confused with this one. Please help. Solve, x½ – 4x¼ = –3
Answer by stanbon(75887)

Solve, x½ – 4x¼ = –3
x^(1/2) - 4x^(1/4) + 3 = 0
Its a quadratic; the variable is x^(1/4)
Factor as follows:
(x^1/4 -3)(x^1/4-1) = 0
x^(1/4) = 3 or x^(1/4)=1
x = 81 o x = 1
======================
Cheers,
Stan H.